A wire of length 2m and radius 2mm is fixed to the center of wheel a torque of magnitude 0.0395Nm is applied to twist the wire find the regidity modulus of wire if angular twist is 0.038radian
1
Expert's answer
2019-10-22T10:44:30-0400

Rigidity modulus of wire


We need to find the rigidity modulus of wire if angular twist is 0.038 radian


Solution:


Length of the wire = l = 2 m


Angular twist = "\theta" = 0.038 radians


Torque magnitude = "{T}" = 0.0395 Nm


Rigidity modulus = "{G}"


Torsion constant =

"{J} =\frac { 1 } {2} \pi r^4"

It depends on the radius of the radius of the wire


Here,

"r = 2 mm = 2 \times 10^{-3} m"



Plug this r in the equation of J


"{J} =\frac { 1 } {2} \pi r^4 =\frac {1}{2} \pi(2 \times 10^{-3})^4 = 2.52 \times 10^{-11} m^4"

We know the formula for Rigidity modulus



"{G} =\frac {T l} {\theta J} = \frac {0.0395 \times 2} {0.038 \times2.52\times 10^{-11}}"


"{G} = 8.2 \times 10^{10} \frac {N} {m^2}"

Answer:

Rigidity modulus =

"{G} = 8.2 \times 10^{10} \frac {N} {m^2}"